Suppose you have Avogadro's number of mini marshmallows and use them to cover the state of Iowa which has a land area of 5.587×10^4 mi^2. Each mini marshmallow has a diameter of 0.635 cm and a height of 2.54 cm. Assuming the marshmallows are packed together so there is no space between them, to what height above the surface, in kilometers, will the mini marshmallows extend?
1
Expert's answer
2019-10-07T06:41:36-0400

1 sq. mile = 2.59 km2

S(Iowa) = 5.587×104*2.59*106=1.4470*1011 m2

h=V/S=(п(d/2)2)*h*N/S=3.14*(0.635*10-2/2)2*2.54*10-2*6.022*1023/1.4470*1011=3.345977*106 m = 3346 km

Answer: 3346 km.